This problem has me stumped. My calculator displays an error too! Could someone please help me?
"Use a graphing calculator to solve.
-5x+8y=7
3y=-2x-9"

Answer by Alan3354(69443)

-5x + 8y = 7
3y = -2x - 9
--------------
2x + 3y = -9 Eqn 1
5x - 8y = -7 Eqn 2
-------
10x + 15y = -45 Eqn 1 times 5
10x - 16y = -14 Eqn 2 times 2
-------------------------------------- Subtract
31y = -31
y = -1
===============
etc
What's the big deal?